what is the difference between the sum of 1/2 and 1/4 and the product of 1/2 and 1/4?

Answer by Edwin McCravy(20055)

"Difference between" means to subtract the first minus the second
"Sum of" means to add the first plus the second
"Product of" means to multiply the first times the second. 


Group the words like this

what is [the difference between (the sum of 1/2 and 1/4) and (the product of 1/2 and 1/4)]?
Then we do (the sum of 1/2 and 1/4) = 1/2 + 1/4 = 2/4 + 1/4 = 3/4.

So we can put 3/4 instead of (the sum of 1/2 and 1/4)

what is [the difference between 3/4 and (the product of 1/2 and 1/4)]?
Then we do (the product of 1/2 and 1/4) = (1/2)(1/4) = 1/8.

So we can put 1/8 instead of (the product of 1/2 and 1/4)

what is [the difference between 3/4 and 1/8]?
Su we just subtract 3/4 - 1/8 = 6/8 - 1/8 = 5/8.

Edwin